OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). -21 -3 3 21

OpenStudy (johnweldon1993):

So we need first find what g(-1) equals. \[\large g(x) = x^2 - 7x - 9\] g(-1) simply means plugging in -1 for every 'x' you see there \[\large g(-1) = (-1)^2 - 7(-1) - 9\] Tell me what that comes out to

hero (hero):

You can also take the approach of finding f(g(x)) first: f(g(x)) = g(x) - 2 = x^2 - 7x - 9 - 2 = x^2 - 7x - 11 Then finding f(g(-1)) from there.

OpenStudy (anonymous):

Okey hold on..

OpenStudy (anonymous):

I got -3

OpenStudy (johnweldon1993):

That is correct!

OpenStudy (anonymous):

Thankyou!